# MAPSS: Manifold-based Assessment of Perceptual Source Separation

### v1

@@ Abstract
Objective assessment of source-separation systems still mismatches subjective human perception, especially when leakage and self-distortion interact. We introduce the Perceptual Separation (PS) and Perceptual Match (PM), the first pair of measures that functionally isolate these two factors. Our intrusive approach begins with generating a bank of fundamental distortions for each reference waveform signal in the mixture. Distortions, references, and their respective system outputs from all sources are then independently encoded by a pre-trained self-supervised learning model. These representations are aggregated and projected onto a manifold via a nonlinear dimensionality reduction technique called diffusion maps, which aligns Euclidean distances on the manifold with dissimilarities of the encoded waveform representations. On this manifold, the PM measures the Mahalanobis distance from each output to its attributed cluster that consists of its reference and distortions embeddings, capturing self-distortion. The PS accounts for the Mahalanobis distance of the output to the attributed and to the closest non-attributed clusters, quantifying leakage. Both measures are differentiable and operate at a resolution as low as 50 frames per second, allowing granular optimization and analysis of self-distortion and leakage for every output channel of the separation system. We further derive, for both measures, deterministic error radius and non-asymptotic, high-probability confidence intervals (CIs). Experiments on English, Spanish, and music mixtures show that compared to 14 alternative measures, the PS and PM nearly always achieve the highest linear correlation coefficients with human mean-opinion scores, reaching as high as 86.36% for speech and 87.21% for music. We observe, at worst, an error radius of $1.39\%$ and a probabilistic 95% CI of 12.21% for these coefficients, which improves reliable and informed evaluation. Using mutual information, we show that the measures strongly complement each other, and more so as their values decrease, suggesting they are jointly more informative as system performance degrades.  [^1] Code available at https://github.com/Amir-Ivry/MAPSS-measures
@@ 1 Introduction
Progress in source-separation heavily depends on reliable perceptual evaluation, but gold-standard human listening tests remain prohibitively costly and time-consuming ITU-T. (1996; 2003; 2018). Consequently, research commonly relies on objective metrics for model development and comparison. However, the most widely adopted metrics conflate two perceptually distinct failure modes: interference from competing talkers and distortion of the target signal. Measures that better isolate these two modes could help close the gap between automatic evaluation and what listeners actually perceive, potentially enabling an accelerated and more trustworthy development of source separation.
@@
Existing measures typically fall into several groups. Signal-to-distortion ratio (SDR), signal-to-interferences ratio (SIR), signal-to-artifacts ratio (SAR) Vincent et al. (2006), scale-invariant SDR (SI-SDR) Le Roux et al. (2019) and alike usually compute ratios between source to various disturbances purely with waveform energy, offering low complexity and widespread adoption. However, even jointly, they cannot tell by design whether an error stems from leakage or self-distortion. Classical intrusive perceptual and intelligibility metrics like the PESQ Rix et al. (2001), STOI Taal et al. (2011) and ESTOI Jensen & Taal (2016) map an entire utterance to a mean-opinion score (MOS) scale using hand-crafted auditory features. Designed preliminary for speech enhancement, they perform well for corrupted noisy-reverberant speech utterances but may not account for leakage, while also lacking to provide access to their inherent granular processing. Learned black-box metrics such as the DNSMOS family Reddy et al. (2022) that are trained end-to-end to predict crowd-sourced MOS, as well as SpeechBERT Tseng et al. (2021) and Sheet-SSQA Huang et al. (2025), have shown promising results on various speech tasks, but do not offer confidence in their decisions. Spectral-distance metrics are interpretable but tend to mask where degradations occur, e.g., the popular Mel-Cepstral Distortion (MCD) Fukada et al. (1992) collapses the spectral envelope into a global value. Even when taking into account a broader set of metrics, e.g., as available in recently developed speech quality assessment toolkits Shi et al. (2024), no existing family of measures can simultaneously disentangle leakage from distortion, offer granular analysis, and provide error estimates for their decisions.
@@
In this paper, we introduce the Perceptual Separation (PS) and Perceptual Match (PM), the first measures for source separation that functionally disentangle leakage and self-distortion. For every reference frame, at a 50 and 10 frames-per-second rate for speech and music mixtures, respectively, we synthesize a bank of fundamental distortions motivated by psycho-acoustic and auditory theory Gabrielsson & Sjögren (1979); Jekosch (2004); Wilson & Fazenda (2014); Bannister et al. (2024). Then, a pretrained self-supervised model, e.g., wav2vec 2.0 Baevski et al. (2020), is used to independently encode the waveform of references, distortions, and system outputs across all sources, producing high-dimensional frame-wise representations. These are aggregated and projected via diffusion maps Coifman & Lafon (2006) onto a low-dimensional manifold, where a key property of the diffusion maps ensures that Euclidean distances on this manifold align with dissimilarities between encoded representations. Harnessing this property and the geometric structures that distortions form around their references on the manifold, then the Mahalanobis distance from an output embedding to its own cluster is used to calculate the PM, whereas PS also accounts for the distance to the nearest foreign cluster. Because these manifolds are learned at tens of milliseconds resolution, low PS pinpoints cross-talk frames, while low PM isolates self-distortion, indicating actionable debugging cues. We derive theoretical deterministic error radius and non-asymptotic, high-probability confidence intervals (CIs) for both measures. These bounds enable frame-level guarantees on the reliability of the measures.
Evaluated on the SEBASS database Kastner & Herre (2022) that includes mixtures of English, Spanish, and music, the PS and PM jointly deliver the top linear correlation coefficients with subjective mean-opinion scores (MOS) in nearly every scenario, compared with $14$ competing widely-used measures. Specifically, coefficients reach as high as $87.21\%$ in music mixtures and $86.36\%$ in English mixtures, with maximal error radius of $1.39\%$ and 95% CI of $12.21\%$. We also achieve the highest ranking-based correlation coefficients in music mixtures with a maximum of $85.54\%$, but scoring $84.69\%$ and $83.41\%$ in English and Spanish places the measures in the top-3 and top-5 places, respectively, with an error radius of 0.4% and a 95% CI of $30.03\%$ at worst. These error margins are imperative to an informed and reliable development of more progressed source-separation systems. Empirically, we also show that the normalized mutual information between the measures never exceeds 0.2, and that as their values decrease it approaches zero. This raises the possibility that when the performance of the system degrades, the simultaneous information from both measures becomes more informative for a reliable assessment.
@@ 2 Problem Formulation
We mark column vectors and matrices with bold symbols and the rest with non-bold symbols.
Consider a trial with index $l$ in which a source separation system with index $q$ performs inference on a single-channel mixture, which includes $N^{l}$ time frames where at least two speakers are active Vincent et al. (2018). Let $\mathcal{F}^{l}$ be the set containing these time frame indices, and assume that in time frame $f\in\mathcal{F}^{l}$} there are $N^{l}_{f}\geq 2$ active speakers and their indices are contained in the set $\mathcal{S}^{l}_{f}$. It therefore holds that:
$\left|\mathcal{F}^{l}\right|$ $=N^{l},$ (1)
$\left|\mathcal{S}^{l}_{f}\right|$ $=N^{l}_{f}.$ (2)
Then, we can model the observed mixture as:
$\mathbf{z}^{l}_{f}=\sum_{i\in\mathcal{S}^{l}_{f}}\mathbf{y}^{l}_{i,f}+\mathbf{v}_{f},$ (3)
where $\mathbf{z}^{l}_{f}=\big{[}z^{l}_{f}(n-L+1),\dots,z^{l}_{f}(n)\big{]}^{T}\in\mathbb{R}^{L}$, i.e., frame $f$ has $L$ samples and starts and ends in the nL+1n-L+1-th and the $n$-th sample of its signal, respectively. All vectors in the section follow the same structure and notation as $\mathbf{z}^{l}_{f}$. Each $\mathbf{y}^{l}_{i,f}\in\mathbb{R}^{L}$ is defined as the reference signal of the $i\text{-th}$} source in frame $f$ and holds the exact source instance that is mixed, potentially including reverberation, distortion, noise, or other interference inherent to its original conditions. The term $\mathbf{v}_{f}$ represents bounded system and environmental interference, assuming statistical independence of the sources. The estimation of $\mathbf{y}^{l}_{i,f}$ is denoted $\hat{\mathbf{y}}^{q,l}_{i,f}\in\mathbb{R}^{L}$.
The first contribution of this paper is the introduction of two new intrusive measures to evaluate source-separation systems. For each source index $i$:
  • The perceptual separation (PS) measure quantifies how well $\hat{\mathbf{y}}^{q,l}_{i,f}$ is perceptually separated from all interfering sources $\{\mathbf{y}^{l}_{j,f}\}_{j\neq i}$.
  • The perceptual match (PM) measure quantifies how closely the estimated source $\hat{\mathbf{y}}^{q,l}_{i,f}$ aligns perceptually with its reference $\mathbf{y}^{l}_{i,f}$.
The second contribution of this paper is the development of theoretical deterministic error radius and non-asymptotic, high-probability confidence intervals for both measures.
3 MAPSS: Manifold-based Assessment of Perceptual Speech Separation
This section discusses the theoretical framework and practical steps we apply to calculate the PM and PS measures. We temporarily assume a fixed speech separation system and time frame and omit the indices $l,q,f$ unless otherwise specified.
3.1 Diffusion Maps: Theoretical Foundations
Diffusion maps provide a low-dimensional manifold representation of a set of high-dimensional points, by learning their geometric and structural relationships Coifman & Lafon (2006). Given a set $\mathcal{X}=\{\mathbf{x}_{0},\mathbf{x}_{1},\dots,\mathbf{x}_{N-1}\}$} of $N$ data points of dimension $M$, an affinity matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$} is calculated using a Gaussian kernel:
@@ 3.1 Diffusion Maps: Theoretical Foundations
where $i,j\in\{0,\ldots,N-1\}$ and $\forall i,j:0\leq K_{i,j}\leq 1$, and:
$\sigma_{\mathbf{K}}^{2}=\operatorname{median}\left\{\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|_{2}^{2}\;\middle|\;i\neq j\right\}.$ (5)
To account for non-uniform sampling density of points, an $\alpha$-normalization replaces $\mathbf{K}$ by $\mathbf{K}^{(\alpha)}$:
@@
where $\alpha\in\left[0,1\right]$ and $v_{i}=\sum_{j=0}^{N-1}\mathbf{K}_{i,j}$. Define the diagonal degree-matrix $\mathbf{D}^{(\alpha)}\in\mathbb{R}^{N\times N}$, given by:
$\mathbf{D}^{(\alpha)}=\operatorname{diag}\left(v^{(\alpha)}_{0},\dots,v^{(\alpha)}_{N-1}\right)$, (7)
with:
$v^{(\alpha)}_{i}=\sum_{j=0}^{N-1}\mathbf{K}^{(\alpha)}_{i,j}$, (8)
where $\forall i:v^{(\alpha)}_{i}>0$ by construction. For readability, we neglect the $\alpha$ notation from now on. The probability transition matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$} on $\mathbf{K}$ can be defined as:
$\mathbf{P}=\mathbf{D}^{-1}\mathbf{K}.$ (9)
Note $\mathbf{P}$ is row-stochastic, so $\forall i,j:\mathbf{P}_{ij}\geq 0,\text{ }\sum_{j=0}^{N-1}\mathbf{P}_{ij}=1$. Spectral decomposition on $\mathbf{P}$ reveals a trivial right eigenvector $\mathbf{u}_{0}=\mathbf{1}\in\mathbb{R}^{N}$} with eigenvalue $\lambda_{0}=1$}. Remaining eigenvectors $\{\mathbf{u}_{\ell}\}_{\ell=1}^{N-1}$ are associated with eigenvalues $\{\lambda_{\ell}\}_{\ell=1}^{N-1}$ ordered as $1>\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N-1}>0$}, so that:
@@
After $t\in\mathbb{Z}^{+}$ time steps on $\mathbf{P}$, the diffusion maps embedding of $\mathbf{x}_{i}$ onto manifold $\mathcal{M}$ is $\Psi_{t}:\mathcal{X}\rightarrow\mathbb{R}^{N-1}$}:
$\mathbf{\Psi}_{t}(\mathbf{x}_{i})=\left(\lambda_{1}^{t}\mathbf{u}_{1}(i),\,\lambda_{2}^{t}\mathbf{u}_{2}(i),\,\dots,\,\lambda_{N-1}^{t}\mathbf{u}_{N-1}(i)\right)^{T},$ (11)
where $\{\mathbf{u}_{\ell}\}_{\ell=1}^{N-1}$ are orthonormal with respect to the stationary measure $\boldsymbol{\pi}=\left[\pi_{0},\ldots,\pi_{N-1}\right]^{T}\in\mathbb{R}^{N\times 1}$} that satisfies:
$\pi_{i}=\frac{\mathbf{D}_{ii}}{\sum_{j=0}^{N-1}\mathbf{D}_{jj}},$ (12)
where $\pi_{i}\in\left(0,1\right)$. Let $D_{t}(i,j)$ be the diffusion distance at time step $t$ between two points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$:
$D^{2}_{t}(i,j)=\sum_{m=0}^{N-1}\frac{\left(\mathbf{P}^{t}_{im}-\mathbf{P}^{t}_{jm}\right)^{2}}{\boldsymbol{\pi}_{m}}$ (13)
where $\mathbf{P}^{t}_{im}$ denotes the probability of transitioning from node $i$ to node $m$ in $t$ time steps. Intuitively, the diffusion distance measures the similarity between the probability distributions of random walks starting from nodes $i$ and $j$. A key strength of diffusion maps we utilize is the following equivalence:
$D^{2}_{t}(i,j)=\big{\|}\boldsymbol{\Psi}_{t}(\mathbf{x}_{i})-\boldsymbol{\Psi}_{t}(\mathbf{x}_{j})\big{\|}_{2}^{2}.$ (14)
This equivalence is fundamental to our approach, since it ensures that Euclidean distances on the manifold align with dissimilarities between their matching high-dimensional points.
In practice, the embedding in (11) is truncated to its first $d$ coordinates, with d<N1d<N-1. This simple yet effective step reduces noise sensitivity and retains only the most meaningful, dominant geometric structures Nadler et al. (2006). Mapping $\Psi_{t}:\mathcal{X}\rightarrow\mathbb{R}^{d}$} onto manifold $\mathcal{M}^{\left(d\right)}$ gives:
$\boldsymbol{\Psi}^{(d)}_{t}(\mathbf{x}_{i})=\left(\lambda_{1}^{t}\mathbf{u}_{1}(i),\,\lambda_{2}^{t}\mathbf{u}_{2}(i),\,\dots,\,\lambda_{d}^{t}\mathbf{u}_{d}(i)\right)^{T}.$ (15)
We cut off the spectral content by choosing $d$ as follows, given $\tau\in\left[0,1\right]$:
$d=\min\left\{k\in\{1,\dots,N-1\}:\frac{\sum_{\ell=1}^{k}\lambda_{\ell}}{\sum_{\ell=1}^{N-1}\lambda_{\ell}}\geq\tau\right\},$ (16)
where we set $\tau=0.99$. Truncating the spectrum to $d$ dimensions breaks the equality in (14), and leads to a truncation error. In App. (A) we derive the expectation and probabilistic tail bound for this truncation error.

### v2

@@ Abstract
Objective assessment of audio source‑separation systems still mismatches subjective human perception, especially when interference from competing talkers and distortion of the target signal interact. We introduce Perceptual Separation (PS) and Perceptual Match (PM), a complementary pair of measures that, by design, isolate these leakage and distortion factors. Our intrusive approach generates a set of fundamental distortions, e.g., clipping, notch filter, and pitch shift from each reference waveform signal in the mixture. Distortions, references, and system outputs from all sources are independently encoded by a pre-trained self-supervised model, then aggregated and embedded with a manifold learning technique called diffusion maps, which aligns Euclidean distances on the manifold with dissimilarities of the encoded waveform representations. On this manifold, PM captures the self‑distortion of a source by measuring distances from its output to its reference and associated distortions, while PS captures leakage by also accounting for distances from the output to non‑attributed references and distortions. Both measures are differentiable and operate at a resolution as high as 75 frames per second, allowing granular optimization and analysis. We further derive, for both measures, frame-level deterministic error radius and non-asymptotic, high-probability confidence intervals. Experiments on English, Spanish, and music mixtures show that, against 14 widely used measures, the PS and PM are almost always placed first or second in linear and rank correlations with subjective human mean-opinion scores [^1] Code available at https://github.com/Amir-Ivry/MAPSS-measures.
@@ 1 Introduction
Reliable perceptual evaluation is critical for source-separation progress, yet gold-standard listening tests are costly and slow (ITU-T., 1996; 2003; 2018). Thus, research relies on objective metrics that blur two distinct failures, interference from competing talkers and target distortion. Disentangling these modes can better align with listener perception and accelerate trustworthy development.
@@
Existing measures such as the signal-to-distortion ratio (SDR), signal-to-interference ratio (SIR), signal-to-artifacts ratio (SAR) (Vincent et al., 2006), scale-invariant SDR (SI-SDR) (Le Roux et al., 2019) and alike usually compute ratios between source to various disturbances in the waveform domain, offering low complexity and widespread adoption. However, even jointly, they are defined with an intrinsic ambiguity to whether an error stems for leakage or self-distortion. Classical intrusive perceptual and intelligibility metrics like the PESQ (Rix et al., 2001), STOI (Taal et al., 2011) and ESTOI (Jensen and Taal, 2016) map an entire utterance to a scaled mean-opinion score (MOS) using hand-crafted auditory features. Designed preliminary for speech enhancement, they perform well for corrupted noisy-reverberant speech utterances but may not account for leakage, while also lacking to provide access to their inherent granular processing. Learned black-box metrics such as the DNSMOS family (Reddy et al., 2022) that are trained end-to-end to predict crowd-sourced MOS, as well as SpeechBERTscore (Saeki et al., 2024) and Sheet-SSQA (Huang et al., 2025), have shown promising results on various speech tasks, but do not offer confidence in their decisions. Spectral-distance metrics are interpretable but tend to mask where degradations occur, e.g., the popular Mel-Cepstral Distortion (MCD) (Fukada et al., 1992) collapses the spectral envelope into a global value. Even when taking into account a broader set of metrics, as available in recently developed speech quality assessment toolkits (Shi et al., 2024), no existing family of measures can simultaneously disentangle leakage from distortion, offer granular analysis, and provide error estimates for their decisions.
@@
We introduce the Perceptual Separation (PS) and Perceptual Match (PM), the first measures for source separation that functionally disentangle leakage and self-distortion. Inspired by auditory theory (Gabrielsson and Sjögren, 1979; Jekosch, 2004; Wilson and Fazenda, 2014; Bannister et al., 2024), we apply a set of fundamental distortions to every reference waveform, intended to create a wide cover of perceptual auditory field around the reference. These distortions range from mildly-intrusive short-tailed reverberations to highly degrading hard clipping. A pretrained self-supervised model, e.g., wav2vec 2.0 (Baevski et al., 2020), is used to independently encode the waveforms of references, distortions, and system outputs across all sources, in a resolution as high as 75 frames-per-second. These representations are aggregated and projected via a manifold learning technique called diffusion maps (Coifman and Lafon, 2006) onto a low-dimensional manifold. A key property of diffusion maps aligns Euclidean distances between points on the manifold with dissimilarities between their encoded representations. On the manifold, PM quantifies self-distortion by measuring how far an output lies from its attributed reference and the distortions, whereas PS quantifies leakage by comparing these distances with the output proximity to non-attributed references and distortions.
Evaluations on the SEBASS database (Kastner and Herre, 2022) with mixtures of English, Spanish, and music, show that compared to 14 widely used measures, PS and PM almost always achieve first- or second-place rankings in both linear and rank correlations with human scores, with the exception of Spanish rank correlations, where they remain within the top third. We derive granular theoretical deterministic error radius and high-probability confidence intervals (CIs) for both measures, enabling frame-level guarantees on the reliability of the measures. In almost all scenarios, the worst-case error radius would not lower the PS and PM rankings. In addition, the normalized mutual information (NMI) (Danon et al., 2005) between the PS and PM values shows that they are highly complementary.
@@ 2 Problem Formulation
Notational remark. Column vectors and matrices are written in bold and other symbols in non-bold.
Consider a source separation system performing inference on an audio mixture (Vincent et al., 2018). In a time frame $f$ that consists of $L$ samples, let $N_{f}\geq 2$ denote the number of active sources and $\mathcal{S}_{f}$ their index set. The observed mixture $\mathbf{z}_{f}\in\mathbb{R}^{L}$ is modeled as:
$\mathbf{z}_{f}=\sum_{i\in\mathcal{S}_{f}}\mathbf{y}_{i,f}+\mathbf{v}_{f},$ (1)
For $i\in\mathcal{S}_{f}$, we denote $\mathbf{y}_{i,f}\in\mathbb{R}^{L}$ the reference signal of the $i\text{-th}$} source in frame $f$, potentially including interference inherent to its original conditions. $\mathbf{v}_{f}$ represents system and environmental interference, assumed statistically independent of the sources. The estimation of $\mathbf{y}_{i,f}$ is denoted $\hat{\mathbf{y}}_{i,f}$.
Given source indices $i,j\in\mathcal{S}_{f}$ in time frame $f$, our goal is to introduce these two measures:
  • The perceptual separation (PS) measure quantifies how well $\hat{\mathbf{y}}_{i,f}$ is perceptually separated from all interfering sources $\{\mathbf{y}_{j,f}\}_{j\neq i}$.
  • The perceptual match (PM) measure quantifies how closely the estimated source $\hat{\mathbf{y}}_{i,f}$ perceptually aligns with its reference $\mathbf{y}_{i,f}$.
3 Diffusion Maps: Theoretical Foundations
[…]
Diffusion maps is a manifold learning method that represents high-dimensional data in a low-dimensional space by capturing geometric and structural relationships (Coifman and Lafon, 2006). Consider the set $\mathcal{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with $\mathbf{x}_{i}\in\mathbb{R}^{M}$ for all $i$, e.g., feature vectors from wav2vec 2.0 (Baevski et al., 2020). An affinity matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$ is calculated between the high-dimensional vectors:
@@ 3 Diffusion Maps: Theoretical Foundations
where $i,j\in\{1,\ldots,N\}$ and $\forall i,j:0\leq K_{i,j}\leq 1$, and $\sigma_{\mathbf{K}}^{2}=\operatorname{median}\left\{\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|_{2}^{2}\;\middle|\;i\neq j\right\}$. To account for non-uniform sampling density of points, an $\alpha$-normalization replaces $\mathbf{K}$ by $\mathbf{K}^{(\alpha)}$:
@@
where $\alpha\in\left[0,1\right]$ and $v_{i}=\sum_{j=1}^{N}\mathbf{K}_{i,j}$. Then, we define the diagonal degree-matrix $\mathbf{D}^{(\alpha)}$, given by $\mathbf{D}^{(\alpha)}=\operatorname{diag}\left(v^{(\alpha)}_{0},\dots,v^{(\alpha)}_{N-1}\right)\in\mathbb{R}^{N\times N}$, where $v^{(\alpha)}_{i}=\sum_{j=1}^{N}\mathbf{K}^{(\alpha)}_{i,j}$ and $\forall i:v^{(\alpha)}_{i}>0$ by construction. We assume $\alpha$ is fixed and for readability we neglect the $\alpha$ notation from now on.
The probability transition operator $\mathbf{P}$ on $\mathbf{K}$ is defined with (3) as:
$\mathbf{P}=\mathbf{D}^{-1}\mathbf{K}\in\mathbb{R}^{N\times N}.$ (4)
Note $\mathbf{P}$ is row-stochastic, so $\forall i,j:\mathbf{P}_{ij}\geq 0,\text{ }\sum_{j=1}^{N}\mathbf{P}_{ij}=1$. Spectral decomposition on $\mathbf{P}$ reveals a trivial right eigenvector $\mathbf{u}_{0}=\mathbf{1}\in\mathbb{R}^{N}$} with eigenvalue $\lambda_{0}=1$}. Remaining eigenvectors $\{\mathbf{u}_{\ell}\}_{\ell=1}^{N-1}$ are associated with eigenvalues $\{\lambda_{\ell}\}_{\ell=1}^{N-1}$ and ordered as $1>\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N-1}>0$}, so that:
@@
Denoting $\mathbf{u}_{j}(i)$ the $i$-th element of the $j$-th eigenvector, then the embedding of $\mathbf{x}_{i}$ onto manifold $\mathcal{M}$ can be expressed with the eigenfunctions in (5), by the embedding operation $\Psi_{t}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{N-1}$}:
$\mathbf{\Psi}_{t}(\mathbf{x}_{i})=\left(\lambda_{1}^{t}\mathbf{u}_{1}(i),\,\lambda_{2}^{t}\mathbf{u}_{2}(i),\,\dots,\,\lambda_{N-1}^{t}\mathbf{u}_{N-1}(i)\right)^{T}.$ (6)
where $t$ is the number of Markov chain steps, controlling the diffusion scale of the embedding. The eigenvalues in $\{\mathbf{u}_{\ell}\}_{\ell=1}^{N-1}$ are orthonormal under the stationary measure $\bm{\pi}=\left[\pi_{1},\pi_{2},\ldots,\pi_{N}\right]^{T}$}:
$\pi_{i}=\frac{\mathbf{D}_{ii}}{\sum_{j=1}^{N}\mathbf{D}_{jj}},\quad\pi_{i}\in\left(0,1\right).$ (7)
Let $D_{t}(i,j)$ be the diffusion distance at time step $t$ between two points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$:
$D^{2}_{t}(i,j)=\sum_{m=1}^{N}\frac{\left(\mathbf{P}^{t}_{im}-\mathbf{P}^{t}_{jm}\right)^{2}}{\pi_{m}}$ (8)
where $\mathbf{P}^{t}_{im}$ (4) denotes the probability of transitioning from node $i$ to node $m$ in $t$ time steps. Intuitively, the diffusion distance measures the similarity between the probability distributions of random walks starting from nodes $i$ and $j$. A key strength of diffusion maps is the equivalence (6):
$D^{2}_{t}(i,j)=\big\|\bm{\Psi}_{t}(\mathbf{x}_{i})-\bm{\Psi}_{t}(\mathbf{x}_{j})\big\|_{2}^{2},$ (9)
[…]
$\bm{\Psi}^{(d)}_{t}(\mathbf{x}_{i})=\left(\lambda_{1}^{t}\mathbf{u}_{1}(i),\,\lambda_{2}^{t}\mathbf{u}_{2}(i),\,\dots,\,\lambda_{d}^{t}\mathbf{u}_{d}(i)\right)^{T}.$ (10)
Consider $\tau\in\left[0,1\right]$ as the minimal normalized retained sum of the eigenvalues, then $d$ is given by:
$d=\min\left\{k\in\{1,\dots,N\}:\frac{\sum_{\ell=1}^{k}\lambda_{\ell}}{\sum_{\ell=1}^{N}\lambda_{\ell}}\geq\tau\right\}.$ (11)
4 The Perceptual Separation and Perceptual Match Measures